Assume that the average firm in C&J Corporation's industry is expected to grow at a constant rate of 7% and that its dividend yield is 6%. C&J is about as risky as the average firm in the industry and just paid a dividend (D0) of $2.5. Analysts expect that the growth rate of dividends will be 50% during the first year (g0,1 = 50%) and 20% during the second year (g1,2 = 20%). After Year 2, dividend growth will be constant at 7%. What is the required rate of return on C&J's stock? What is the estimated intrinsic price per share? Do not round intermediate calculations. Round the monetary value to the nearest cent and percentage value to the nearest whole number.

Please show work!

Thanks
Answer #1

Required Rate =Dividend Yield +Capital Gain =6%+7% =13%
D1 =D0*(1+20%) =2.5*(1+20%) =3
D2=D0*(1+20%)^2 =2.5*1.20^2 =3.60
D3=D2*(1+growth) =3.60*(1+7%) =3.852
Terminal Value =D3/(Required Rate-Growth)=3.852/(13%-7%) =64.20

Estimated Intrinsic Value =D1/(1+r)+D2/(1+r)^2+Terminal value/(1+r)^2
=3/(1+13%)+3.60/(1+13%)^2+64.20/(1+13%)^2 =55.75